# Learning Adaptive and View-Invariant Vision Transformer with Multi-Teacher Knowledge Distillation for Real-Time UAV Tracking

### v1

@@ 4.5 Qualitative Results:
Visualization of feature maps: In Fig. 6, the first row presents the original target images from UAV123@10fps (Mueller et al., 2016), while the corresponding feature maps generated by AVTrack-DeiT*, AVTrack-DeiT, and AVTrack-MD-DeiT are displayed in the second, third, and fourth rows, respectively. Note that AVTrack-DeiT* refers to AVTrack-DeiT without the integration of the proposed VIR and AM components. As observed, the feature maps generated by AVTrack-DeiT and AVTrack-MD-DeiT demonstrate greater consistency with changes in viewpoint, while the feature maps from AVTrack-DeiT* exhibit more significant variations, particularly at different viewing angles. This suggests that the VIR component enhances the tracker’s ability to focus on targets experiencing changes in viewpoint, thereby improving overall tracking performance. Furthermore, the proposed MD allows AVTrack-MD-DeiT to inherit this capability from AVTrack-DeiT, enabling it to maintain similar robustness to the teacher models while also enhancing model efficiency. These qualitative results offer visual evidence of our method’s effectiveness in learning view-invariant feature representations using Vision Transformers (ViTs).
@@
4.6 Real-world Test
To showcase the applicability of AVTrack-MD-DeiT for UAV tracking, we performed two real-world tests on the NVIDIA Jetson AGX Xavier 32GB. As shown in Fig. 7, the proposed tracker performs well in real-time despite challenges like camera motion, occlusion, and viewpoint changes, maintaining a Center Location Error (CLE) of under 20 pixels across all test frames, highlighting its high tracking precision. Furthermore, AVTrack-MD-DeiT consistently achieves an average real-time speed of 46 FPS in tests, demonstrating its robustness and efficiency, making it ideal for UAV applications requiring high performance and speed.
5 Conclusions
In this paper, we explore the effectiveness of a unified framework (AVTrack) for real-time UAV tracking through the use of efficient Vision Transformers (ViTs). To achieve this, we introduced an adaptive computation paradigm that selectively activates Transformer blocks. Additionally, to address the challenges of varying viewing angles commonly faced in UAV tracking, we utilized mutual information maximization to learn view-invariant representations. Building on this, we have developed an improved AVTrack-MD model to enable more efficient UAV tracking by introducing an simple and effective MI maximization-based multi-teacher knowledge distillation (MD) framework. Extensive experiments across six UAV tracking benchmarks show that AVTrack-MD not only delivers performance comparable to the AVTrack but also reduces model complexity, leading to a substantial improvement in tracking speed.
Acknowledgements
Thanks to the support by the Guangxi Natural Science Foundation (Grant No. 2024GXNSFAA010484), and the National Natural Science Foundation of China (Nos. 62466013, 62206123, 62066042, and 62176170), this work has been made possible.

### v2

@@ IV-E Qualitative Results:
Visualization of feature maps: In Fig. 6, the first row presents the original target images from UAV123@10fps [88], while the corresponding feature maps generated by AVTrack-DeiT*, AVTrack-DeiT, and AVTrack-MD-DeiT are displayed in the second, third, and fourth rows, respectively. Note that AVTrack-DeiT* refers to AVTrack-DeiT without the integration of the proposed VIR and AM components. As observed, the feature maps generated by AVTrack-DeiT and AVTrack-MD-DeiT demonstrate greater consistency with changes in viewpoint, while the feature maps from AVTrack-DeiT* exhibit more significant variations, particularly at different viewing angles. This suggests that the VIR component enhances the tracker’s ability to focus on targets experiencing changes in viewpoint, thereby improving overall tracking performance. Furthermore, the proposed MD allows AVTrack-MD-DeiT to inherit this capability from AVTrack-DeiT, enabling it to maintain similar robustness to the teacher models while also enhancing model efficiency. These qualitative results offer visual evidence of our method’s effectiveness in learning view-invariant feature representations using Vision Transformers (ViTs).
@@
IV-F Real-world Test
To showcase the applicability of AVTrack-MD-DeiT for UAV tracking, we performed three real-world tests on the NVIDIA Jetson AGX Xavier 32GB. As shown in Fig. 7, the proposed tracker performs well in real-time despite challenges like camera motion, occlusion, and viewpoint changes, maintaining a Center Location Error (CLE) of under 20 pixels across all test frames, highlighting its high tracking precision. Furthermore, AVTrack-MD-DeiT consistently achieves an average real-time speed of 46 FPS in tests, demonstrating its robustness and efficiency, making it ideal for UAV applications requiring high performance and speed.
V Conclusions
In this paper, we explore the effectiveness of a unified framework (AVTrack) for real-time UAV tracking through the use of efficient Vision Transformers (ViTs). To achieve this, we introduced an adaptive computation paradigm that selectively activates Transformer blocks. Additionally, to address the challenges of varying viewing angles commonly faced in UAV tracking, we utilized mutual information maximization to learn view-invariant representations. Building on this, we have developed an improved AVTrack-MD model to enable more efficient UAV tracking by introducing a simple and effective MI maximization-based multi-teacher knowledge distillation (MD) framework. Extensive experiments across six UAV tracking benchmarks show that AVTrack-MD not only delivers performance comparable to the AVTrack but also reduces model complexity, leading to a substantial improvement in tracking speed.